# Probing Non-Fermi-Liquid Behaviour of Composite Fermi Liquid via Efficient Thermal Simulations

### v1

@@ Abstract
We can also demonstrate the formation of the CFL state as a function of temperature. In the 2D limit, the composite Fermion in CFL are expected to form a circular Fermi surface with Fermi momentum $k_{F}=1$. But it is not manifested in the spectral function of electrons, rather, due to the non-Fermi-liquid nature of the state, can be verified by the “guiding center” density-density correlation $\bar{D}(q)=\exp(q^{2}/2)\langle n_{q}n_{-q}\rangle_{T}=\exp(q^{2}/2)\operatorname{Tr}[n_{q}n_{-q}\exp(-\mathcal{H}/T)]/Z$, which reflects the scattering events of CFs near the composite Fermi surface with a $2k_{F}$ singularity circle in the momentum space. In Fig. 4 (a), we plot $\bar{D}(q)$ as a function of 2D momenta $q=(q_{x},q_{y})$ at four different temperatures $T=10^{1},10^{0},10^{-1},10^{-2}$. It starts to exhibit a $2k_{F}$ circle for $T\lesssim 10^{-1}$, i.e. the evidence for the formation of the composite Fermi surface. As shown in the panel (b) and (c), we choose two fixed-$q_{y}$ (specifically, $q_{y}=0,2\pi/L_{y}$) cuts in the 2D momentum contour. The dashed grey lines indicate the two $2k_{F}$ points for the corresponding $q_{y}$’s. One can then see a clear round peak around the $2k_{F}$ points, which will become the true singularity points in the 2D limit, and the round peaks build up significantly around $T\sim 10^{-1}$ manifesting our finding above. This can also be seen in the panel (d), where $\bar{D}(q)$ at the three different $2k_{F}$ points of $q_{y}=m\frac{2\pi}{L_{y}}$ with $m=0,1,2$, which increase most quickly at $\sim 10^{-1}$. Such temperature scale matches with that of non-Fermi-liquid specific heat observed in Fig. 3. We note that, in the low-temperature regime, the small-$q$ density correlation $D(q)$ shows a Fermi-liquid $q^{3}$ behaviour [c.f. SM [83]]. This implies the CFs by themselves, to some extent, behave as a Fermi liquid, and it is the strong coupling with the emergent gauge field that contributes to the non-Fermi-liquid behaviour of the specific heat.
@@
Furthermore, our findings demonstrate the efficacy of the thermal tensor network approach within the projected Hamiltonian framework, particularly in momentum space, highlighting its potential utility in continuum models in the context of two-dimensional moiré materials to study the magnetism and topology [89, 90, 91].
Acknowledgment.- We thank Dung Xuan Nguyen, Bo Yang and Duncan Haldane for discussion on the subject. We acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17309822, C7037-22GF, 17302223, 17301924), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services at the University of Hong Kong, as well as the Beijing Paratera Tech Corp., Ltd [92] for providing HPC resources that have contributed to the research results reported within this paper.

### v2

@@ Abstract
We can also demonstrate the formation of the CFL state during the cooling process. In the 2D limit, the composite Fermion in CFL are expected to form a circular Fermi surface with Fermi momentum $k_{F}=1$. But it is not manifested in the spectral function of the bare electrons, rather, due to the non-Fermi-liquid nature of the state, can be verified by the “guiding center” density-density correlation $\bar{D}(q)=\exp(q^{2}/2)\langle n_{q}n_{-q}\rangle_{T}=\exp(q^{2}/2)\operatorname{Tr}[n_{q}n_{-q}\exp(-\mathcal{H}/T)]/Z$, which reflects the scattering events of CFs near the composite Fermi surface with a $2k_{F}$ singularity circle in the momentum space. In Fig. 4 (a), we plot $\bar{D}(q)$ as a function of 2D momenta $q=(q_{x},q_{y})$ at four different temperatures $T=10^{1},10^{0},10^{-1},10^{-2}$. It starts to exhibit a $2k_{F}$ circle for $T\lesssim 10^{-1}$, i.e. the evidence for the formation of the composite Fermi surface. As shown in the panel (b) and (c), we choose two fixed-$q_{y}$ (specifically, $q_{y}=0,2\pi/L_{y}$) cuts in the 2D momentum contour. The dashed grey lines indicate the $2k_{F}$ points for the corresponding $q_{y}$’s. One can then see a clear round peak around the $2k_{F}$ points, which will become the true singularity points in the 2D limit, and the round peaks build up significantly around $T\sim 10^{-1}$ manifesting our finding above. This can also be seen in the panel (d), where $\bar{D}(q)$ at the three different $2k_{F}$ points of $q_{y}=m\frac{2\pi}{L_{y}}$ with $m=0,1,2$, which increase most quickly at $\sim 10^{-1}$. Such temperature scale matches with that of non-Fermi-liquid specific heat observed in Fig. 3. We note that, in the low-temperature regime, the small-$q$ density correlation $D(q)$ shows a Fermi-liquid $q^{3}$ behaviour [c.f. SM [95]]. This implies the CFs by themselves, to some extent, behave as electrons in conventional Fermi liquid, and it is the strong coupling with the emergent gauge field that contributes to the non-Fermi-liquid behaviour of the specific heat.
@@
Furthermore, our findings demonstrate the efficacy of the thermal tensor network approach within the framework of projected Hamiltonian [91, 92], particularly in momentum space, highlighting its potential utility in continuum models in the context of other FQH systems [96] and FQAH and quantum moiré materials – only need to change the basis function – to study the thermodynamics, magnetism and topology [97, 98, 99]. E.g., our methodological can be straightforwardly applied to the zero-field CFL in the twisted TMD system [61, 62, 57, 60] to explore whether such a sublinear-$T$ specific heat still holds in the absence of Landau level, which might offer a thermal signature that distinguish the CFL from Fermi liquid in the quantum moiré experiments.
Acknowledgment.- We thank Dung Xuan Nguyen, Bo Yang and Duncan Haldane for discussion on the subject. We acknowledge the support from the Research Grants Council (RGC) of Hong Kong (Project Nos. 17309822, C7037-22GF, 17302223, 17301924, 17301725), the ANR/RGC Joint Research Scheme sponsored by RGC of Hong Kong and French National Research Agency (Project No. A_HKU703/22). We thank HPC2021 system under the Information Technology Services at the University of Hong Kong [100], as well as the Beijing Paratera Tech Corp., Ltd [101] for providing HPC resources that have contributed to the research results reported within this paper.